OpenStudy (anonymous):

The distance d that an object falls is directly proportional of the square of the time of the fall, t. A person who is parachuting for the first time is told to wait 19 seconds before opening the parachute. If the person falls 64 feet in 2 seconds, find how far he falls in 19 seconds.

OpenStudy (anonymous):

64/2 = 32 32x 19 = 608

OpenStudy (anonymous):

I believe this is your final answer if all constants are fixed and no gravity or other factors are applied.

OpenStudy (anonymous):

ok thanks man

OpenStudy (anonymous):

d = kt^2 64 = k*2^2 k = 16 so d = 16t^2 when t = 19 d = 16*19^2 d = 5776

OpenStudy (anonymous):

oops he is right ^ I guess I didn't read the question fully.

OpenStudy (anonymous):

yea - its t^2 - thats what you missed

OpenStudy (anonymous):

oh alright thanks man